OpenStudy (dmezzullo):

What is the slope of the line 2x + 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (dmezzullo):

@nubeer

OpenStudy (amistre64):

make x=0, what is y? make x=1, what is y? what is the change in y from x=0 to x=1?

OpenStudy (dmezzullo):

1 is the change

OpenStudy (amistre64):

no, it isnt

OpenStudy (dmezzullo):

and idk bout the 1st 2 things

OpenStudy (dmezzullo):

wut is it then?

OpenStudy (zpupster):

Ax + By = C y = -A/B x + C/B -A/B is the slope

OpenStudy (amistre64):

slope is the change in the value of y, as x changes by 1

OpenStudy (amistre64):

2(0) + 5y = -10 , solve for y 2(1) + 5y = -10 , solve for y what is the change in y, from x=0 to x=1?

OpenStudy (anonymous):

uhhhm i would have to say c... i think..

OpenStudy (anonymous):

oops i mean -2/5

OpenStudy (dmezzullo):

(1) is the change?

OpenStudy (amistre64):

when x=0 2(0) + 5y = -10 , solve for y

OpenStudy (dmezzullo):

-5y on both

OpenStudy (dmezzullo):

sides

OpenStudy (amistre64):

2(0) + 5y = -10 5y = -10 5(-2) = -10 .... when x=0, y = -2

OpenStudy (dmezzullo):

ohhh

OpenStudy (amistre64):

when x=1 2(1) + 5y = -10 2 + 5y = -10 5y = -12 when x=1, y= -12/5

OpenStudy (amistre64):

the change in y is therefore: \[-\frac{12}5-(-2)\] \[-\frac{12}5+\frac{10}5\] \[\frac{-12+10}5\] \[\frac{-2}5\]

OpenStudy (anonymous):

yays i got it riight

OpenStudy (dmezzullo):

Okie dokie thanks alot

OpenStudy (amistre64):

it is important to know what slope IS, and then you can be able to work it into properties such as: ax+by = c, slope is -a/b

OpenStudy (dmezzullo):

Oh i see